# PaperArena: An Evaluation Benchmark for Tool-Augmented Agentic Reasoning on Scientific Literature

### v1

@@ 1. Introduction
Despite this potential, progress in LLM-based agents for scientific research remains limited. A primary obstacle is the lack of a benchmark and an evaluation platform designed to assess agents’ ability to address complex real-world research questions, which usually involve cross-document reasoning and coordinated tool usage. As shown in Table 1, existing benchmarks such as PubMedQA (Jin et al., 2019) and SPIQA (Pramanick et al., 2024) focus on question answering (QA) over scientific documents, but they typically restrict tasks to single-paragraph, single-modality settings, such as explaining a concept from a paragraph or extracting a value from a figure. However, modern advanced LLMs can often solve these directly using their inherent capabilities without tool invocation, offering limited challenge to tool-augmented agents. Meanwhile, although recent efforts such as ReAct (Yao et al., 2023) and ToolLLM (Qin et al., 2023) begin to explore the tool-use capabilities of agents, they lack the specific adaptations of navigating and reasoning on scientific literature.
@@ 2.1. Benchmarks on Scientific Literature
Benchmarks for scientific literature understanding have rapidly grown in the past few years.  (Hendrycks et al., 2020; Saikh et al., 2022). Early benchmarks target specific tasks such as claim verification (Wadden et al., 2020) and phrases extraction (Augenstein et al., 2017). With the rise of LLMs, inquiries on knowledge-intensive scientific data receive increasing attention. For example, SciQA (Auer et al., 2023) focuses on scientific QA over structured academic knowledge. In specialized domains, PubMedQA (Jin et al., 2019) and BioASQ (Krithara et al., 2023) focus on biomedical literature, while ChemLLMBench (Guo et al., 2023) extends evaluation to chemical challenges. Recently, LLM-based agents show promising potential, inspiring the development of expert-level benchmarks for scientific literature mining. In terms of reproducibility, ResearchCodeBench (Hua et al., 2025) evaluates machine learning code completion, whereas PaperBench (Starace et al., 2025) centers on reproducing code from scratch, providing only a paper. Numerous task-oriented benchmarks are also proposed to evaluate agents (Majumder et al., 2024; Chen et al., 2024; Zhou et al., 2025; Cui et al., 2025; Laurent et al., 2024). In the domain of scientific inquiry, PeerQA (Baumgärtner et al., 2025) compiles QA pairs from peer review questions, and Humanity’s Last Exam (Phan et al., 2025) assesses LLM performance in high-difficulty professional questions.
@@ 2.2. Tool-Augmented Agents and Evaluation
The rapid advancement of LLMs and their emerging reasoning capabilities (Brown et al., 2020; Wei et al., 2022; Achiam et al., 2023) pave the way for the development of agent workflows. Early agents demonstrated the extension of reasoning through external interactions. WebGPT (Nakano et al., 2021) showed LLMs retrieving and synthesizing information via web browsing, while SayCan (Ahn et al., 2022) highlighted planning and acting in embodied tasks. Building on these ideas, ReAct (Yao et al., 2023) introduced a paradigm that synergizes reasoning and actions, marking the beginning of a new era for tool-augmented reasoning agents. Following ReAct, research has expanded the capabilities of agents in planning (Huang et al., 2024), tool use (Qiu et al., 2025), memory (Zhang et al., 2024; Li et al., 2025), and reflection (Shinn et al., 2023). Recently, there are also some influential works in multi-agent systems (Hong et al., 2024; Wu et al., 2024; Zhang et al., 2025a; Ye et al., 2025; Wang et al., 2025b; Qiao et al., 2025). From the perspective of evaluation (Wang et al., 2025a), PaperQA (Lála et al., 2023; Skarlinski et al., 2024) evaluate agents on real-world literature retrieval tasks and STELLA (Jin et al., 2025) evaluate agents through a dynamic tool ocean to advance biological research. However, existing works are restricted to tool-free tasks within isolated papers, with no effective benchmarks for tool-centric complex literature understanding. To bridge this gap, we propose PaperArena, a new benchmark for tool-augmented agents on scientific literature, and its corresponding evaluation platform.
@@ 4.3. Benchmark Evaluation
We evaluate nine state-of-the-art LLMs in both single-agent and multi-agent systems: Gemini 2.5 Pro (Comanici et al., 2025), OpenAI o4-mini-high (OpenAI, 2025b), Claude Sonnet 4 (Anthropic, 2025), GPT-4.1 (OpenAI, 2025a), Claude 3.5 Sonnet (Anthropic, 2024), Qwen3-235B-Thinking (Yang et al., 2025), GLM-4.5 (Zeng et al., 2025), Qwen3-235B-Instruct (Yang et al., 2025), and Kimi-K2-Instruct (Team K et al., 2025). We construct a 200-question subset from PaperArena (PaperArena-Human) and recruit three computer science Ph.D. candidates to complete it using standard tools such as GPT-4o and web search engines. All answers both from LLMs and human experts are evaluated using an LLM-as-a-Judge protocol, where GPT-4o assigns binary correctness labels. To emphasize reasoning over surface form, we allow flexible output styles as long as the reasoning is valid and the final answer is unambiguous. In addition to correctness, we measure agent behavior using two metrics: (1) average reasoning steps, defined as the mean length of executed tool chains; and (2) average reasoning efficiency, computed as the average overlap between the agent’s executed tool chain and the predefined chain, normalized by the executed length. To ensure evaluation reliability, we manually verify all human answers and a sample of LLM responses within the human subset.
@@ Appendix A Ethical Statement
We strictly adhere to established ethical standards throughout the benchmark construction process. All scientific papers included in this work were retrieved exclusively from publicly accessible open-access platforms such as OpenReview. The sampling relied entirely on LLMs without manual preference or bias. These papers were collected solely for non-commercial academic research purposes. We explicitly acknowledge the copyright of the original authors and confirm that our use of the papers does not infringe on proprietary or sensitive content. All generated questions, metadata, and annotations are released under an open license, and we commit to making the dataset and associated code publicly available. This benchmark is intended solely for scientific research, and we explicitly prohibit any commercial use or redistribution of the underlying data.
Users of our benchmark must also comply with strict ethical standards. The dataset is released under a non-commercial license for academic research only. Any form of commercial deployment, redistribution, or monetization of the benchmark, tools, or derived models is strictly forbidden. Users are required to provide proper attribution by citing this work when using any part of the benchmark and to ensure downstream applications comply with data usage restrictions and licensing terms set by the original paper sources.
@@ C.3. Analysis of Reasoning Errors
Context Retriever Performs semantic search over long texts to find relevant passages. Query string. Relevant text chunks. query, top_k Uses text-embedding-3-small to embed the query and chunks, then performs similarity search.

### v2

@@ 1. Introduction
Despite this potential, progress in LLM-based agents for scientific research remains limited. A primary obstacle is the lack of a benchmark and an evaluation platform designed to assess agents’ ability to address complex real-world research questions, which usually involve cross-document reasoning and coordinated tool usage. As shown in Table 1, existing benchmarks such as PubMedQA (Jin et al., 2019) and SPIQA (Pramanick et al., 2024) focus on question answering (QA) over scientific documents, but they typically restrict tasks to single-paragraph, single-modality settings, such as explaining a concept from a paragraph or extracting a value from a figure. However, modern advanced LLMs can often solve these directly using their inherent capabilities without tool invocation, offering limited challenge to tool-augmented agents (Zhou et al., 2025; Li et al., 2025a; Cheng et al., 2025). Meanwhile, although recent efforts such as ReAct (Yao et al., 2023) and ToolLLM (Qin et al., 2023) begin to explore the tool-use capabilities of agents, they lack the specific adaptations of navigating and reasoning on scientific literature.
@@ 2.1. Benchmarks on Scientific Literature
Benchmarks for scientific literature understanding have rapidly grown in the past few years.  (Hendrycks et al., 2020; Saikh et al., 2022). Early benchmarks target specific tasks such as claim verification (Wadden et al., 2020) and phrases extraction (Augenstein et al., 2017). With the rise of LLMs, inquiries on knowledge-intensive scientific data receive increasing attention. For example, SciQA (Auer et al., 2023) focuses on scientific QA over structured academic knowledge. In specialized domains, PubMedQA (Jin et al., 2019) and BioASQ (Krithara et al., 2023) focus on biomedical literature, while ChemLLMBench (Guo et al., 2023) extends evaluation to chemical challenges. Recently, LLM-based agents show promising potential, inspiring the development of expert-level benchmarks for scientific literature mining. In terms of reproducibility, ResearchCodeBench (Hua et al., 2025) evaluates machine learning code completion, whereas PaperBench (Starace et al., 2025) centers on reproducing code from scratch, providing only a paper. Numerous task-oriented benchmarks are also proposed to evaluate agents (Majumder et al., 2024; Chen et al., 2024; Cui et al., 2025; Laurent et al., 2024; Ouyang et al., 2025). In the domain of scientific inquiry, PeerQA (Baumgärtner et al., 2025) compiles QA pairs from peer review questions, and Humanity’s Last Exam (Phan et al., 2025) assesses LLM performance in high-difficulty professional questions.
@@ 2.2. Tool-Augmented Agents and Evaluation
The rapid advancement of LLMs and their emerging reasoning capabilities (Brown et al., 2020; Wei et al., 2022; Achiam et al., 2023; Li et al., 2025a; Jiang et al., 2025) pave the way for the development of agent workflows. Early agents demonstrated the extension of reasoning through external interactions. WebGPT (Nakano et al., 2021) showed LLMs retrieving and synthesizing information via web browsing, while SayCan (Ahn et al., 2022) highlighted planning and acting in embodied tasks. Building on these ideas, ReAct (Yao et al., 2023) introduced a paradigm that synergizes reasoning and actions, marking the beginning of a new era for tool-augmented reasoning agents. Following ReAct, research has expanded the capabilities of agents in planning (Huang et al., 2024), tool use (Qiu et al., 2025), memory (Zhang et al., 2024; Li et al., 2025b), and reflection (Shinn et al., 2023). Recently, there are also some influential works in multi-agent systems (Hong et al., 2024; Wu et al., 2024; Zhang et al., 2025a; Ye et al., 2025; Wang et al., 2025b; Qiao et al., 2025). From the perspective of evaluation (Wang et al., 2025a), PaperQA (Lála et al., 2023; Skarlinski et al., 2024) evaluate agents on real-world literature retrieval tasks and STELLA (Jin et al., 2025) evaluate agents through a dynamic tool ocean to advance biological research. However, existing works are restricted to tool-free tasks within isolated papers, with no effective benchmarks for tool-centric complex literature understanding. To bridge this gap, we propose PaperArena, a new benchmark for tool-augmented agents on scientific literature, and its corresponding evaluation platform.
@@ 4.3. Benchmark Evaluation
We evaluate nine state-of-the-art LLMs in both single-agent and multi-agent systems: Gemini 2.5 Pro (Comanici et al., 2025), OpenAI o4-mini-high (OpenAI, 2025b), Claude Sonnet 4 (Anthropic, 2025), GPT-4.1 (OpenAI, 2025a), Claude 3.5 Sonnet (Anthropic, 2024), Qwen3-235B-Thinking (Yang et al., 2025), GLM-4.5 (Zeng et al., 2025), Qwen3-235B-Instruct (Yang et al., 2025), and Kimi-K2-Instruct (Team K et al., 2025). We construct a 200-question subset from PaperArena (PaperArena-Human) and recruit three computer science Ph.D. candidates to complete it using standard tools such as GPT-4o and web search engines. All answers both from LLMs and human experts are evaluated using an LLM-as-a-Judge protocol (Liu et al., 2025), where GPT-4o assigns binary correctness labels. To emphasize reasoning over surface form, we allow flexible output styles as long as the reasoning is valid and the final answer is unambiguous. In addition to correctness, we measure agent behavior using two metrics: (1) average reasoning steps, defined as the mean length of executed tool chains; and (2) average reasoning efficiency, computed as the average overlap between the agent’s executed tool chain and the predefined chain, normalized by the executed length. To ensure evaluation reliability, we manually verify all human answers and a sample of LLM responses within the human subset.
@@ Appendix A Ethical Statement
All source papers were retrieved from public, open-access platforms using unbiased, automated sampling. The dataset is intended exclusively for non-commercial academic research, and all generated artifacts are publicly released under an open license. The benchmark is distributed under a non-commercial license. Users must cite this work and comply with the original papers’ licensing terms. Any commercial use is strictly prohibited.
@@ C.3. Analysis of Reasoning Errors
Context Retriever Performs semantic search over long texts to find relevant passages (Yu et al., 2024). Query string. Relevant text chunks. query, top_k Uses text-embedding-3-small to embed the query and chunks, then performs similarity search.